OpenStudy (anonymous):

Find the foci of x^2 / 9 + y^2 / 36 = 1

OpenStudy (anonymous):

This is an equation of an ellipse with center at the origin (0,0). With \[x^2/(3)^2+ y^2/(6)^2 =1\] Setting a=6, b=3 the foci are at \[(0,\pm c)\]where \[c^2=a^2-b^2\] so \[c=\sqrt{36-9}=\sqrt{25}=5\] The foci are at (0,5) and (0,-5)